A wall has two layers A and B each made of different materials. the layer A is 10 cm thick and B is 20 cm thick. the thermal conductivity of A is thirce that of B. Under thermal equilibrium, temperature across the wall is 35 degree celcius. The difference of temperature across the layer A is
a) 30 degree celcius.
b) 14 degree celcius.
c) 8.75 degree celcius.
d) 5 degree celcius.

Grade:11

2 Answers

Swapnil Saxena
12 years ago

If the walls are in series then the heat currents established in both thew walls should be equal.

Let the temp. at A=0(say) , B=35( to maintain the temp. difference of 35 b/w the walls) and T (say across the layer)

= dQ1/dt =dQ2/dt

= (3K)A(T-0) / 10 = (K)A(35-T) / 20

= (6K)A(T)= (K)A(35-T)

= (6K) = (35-T)

=>7T=35

=> T= 5 -----------(ans)

ankit singh
3 years ago

let the temp. at A=0(say) , B=35( to maintain the temp. difference of 35 b/w the walls) and T (say across the layer)

= dQ1/dt =dQ2/dt

= (3K)A(T-0) / 10 = (K)A(35-T) / 20

= (6K)A(T)= (K)A(35-T)

= (6K) = (35-T)

=>7T=35

=> T= 5 -----------(ans)